tyresseparris30:

Given the system of equations, what is the solution? 2x + y = -1 x - y = -5

563blackghost:

You need to change one of the equation into slope-intercept form. Let's change the second one. \(\large\bf{x-y=-5}\) You would `subtract x` from both sides. \(\bf{x\color{red}{-x}-y=-5\color{red}{-x}}\) \(\bf{-y=-x-5}\) Now rid of the negative. \(\large\bf{y=x+5}\) ~~~ Now you would apply the equation into the first equation. \(\large\bf{2x+(x+5)=-1}\) How would you simplify this?